OpenStudy (anonymous):

Perform the requested operation or operations. f(x) =x-3/5 ; g(x) = 5x + 3, find g(f(x)).

OpenStudy (campbell_st):

all this means is instead of x in g(x) you need to replace it with x - 3/5 so you are looking at \[g(\frac{x - 3}{5}) = 5(\frac{x -3}{5}) + 3\] just distribute and collect liek terms in g(x) hope this helps

OpenStudy (jdoe0001):

$$ f(x) =\color{red}{\frac{x-3}{5}}\\ g(x) = 5x + 3\\ g(f(x)) = 5\pmatrix{\color{red}{\frac{x-3}{5}}} + 3 $$

OpenStudy (anonymous):

=3/5

OpenStudy (anonymous):

f(g(x))=x-3/5

OpenStudy (campbell_st):

not quite distribute and you get \[g(\frac{x -3}{5}) = 5 \times \frac{1}{5} (x - 3) + 3\] so you have \[g(\frac{x -3}{5}) = x - 3 + 3\] just finish it off

OpenStudy (anonymous):

f(g(x))=x+6

OpenStudy (campbell_st):

nope just be careful with the signs -3 + 3 = 0 so what are you left with..?

OpenStudy (jdoe0001):

$$ f(x) =\frac{x-3}{5}\\ g(x) = \color{red}{5x + 3}\\ f(g(x)) = \frac{(\color{red}{5x + 3})-3}{5} $$

OpenStudy (anonymous):

Oh, poop...Thank you so much! I always forget the signs

OpenStudy (campbell_st):

glad to help